# Limit Computation Over Posets via Minimal Initial Functors

### v1

@@ 2.2.1. Limits as the Solutions to Linear Systems
For any small category $C$ and functor $G\colon C\to\mathbf{Vec}$, $\lim G$ admits a simple, concrete description as the solution of a system of linear equations: For $v\in\prod_{c\in\operatorname{Ob}C}G_{c}$ and $b\in\operatorname{Ob}C$, let $v_{b}$ be the projection of $v$ on $G_{b}$. For $\gamma\colon b\to c\in\hom C$, write $s(\gamma)=b$ and $t(\gamma)=c$. A standard result [mac2013categories, Theorem V.2.2] is that
@@
with each cone map $\lim G\to G_{c}$ the restriction of the projection
$\prod_{c\in\operatorname{Ob}C}G_{c}\twoheadrightarrow G_{c}.$
@@
(2.2) $\lim G=\left\{v\in\!\prod_{m\in M}G_{m}\,:\,G_{mq}(v_{m})=G_{m^{\prime}q}(v_{m^{\prime}})\ \forall\,m,m^{\prime}\in M\textup{ with }m,m^{\prime}<q\right\},$
@@
$\lim G\hookrightarrow\prod_{m\in M}G_{m}\twoheadrightarrow G_{m^{\prime}}\xrightarrow{G_{m^{\prime}q}}G_{q},$
for any choice of $m^{\prime}\in M$ with $m^{\prime}\leq q$.
@@
Sketch of proof of Proposition˜2.6.
@@ 2.3. Initial Functors
  • objects are the pairs $(c,\alpha)$, where $c\in\operatorname{Ob}C$ and $\alpha\in\hom(F(c),d)$,
  • morphisms $(c,\alpha)\to(c^{\prime},\alpha^{\prime})$ are morphisms $\kappa\colon c\to c^{\prime}$ in $C$ such that $\alpha=\alpha^{\prime}\circ F(\kappa)$.
@@
The proof of Proposition˜2.10 in [mac2013categories] gives a concrete construction of the cone $\lim G$ from the cone $\lim F\circ G$, where each cone map is taken to be the composition of a cone map of $\lim F\circ G$ with a structure map of $G$. This construction makes precise the idea that to compute $\lim G$, it suffices to compute $\lim G\circ F$.
We will be particularly interested in functors $F\colon C\to Q$ where $Q$ is a poset. In this case, for $q\in Q$, $(F\downarrow q)$ can be identified with the full subcategory of $C$ with objects $\{c\in\operatorname{Ob}C\mid F(c)\leq q\}$. In particular, we have the following:
@@
$\{(m_{1},e_{1}),(m_{2},e_{1}),(m_{2},e_{2}),(m_{3},e_{2}),(e_{1},e_{3}),(e_{2},e_{3}),(m_{4},e_{3})\}.$
For each $i\in\{1,2,3\}$, let $P_{i}\subseteq Q$ be the subposet with $\operatorname{Ob}P_{i}=\operatorname{Ob}Q$, whose Hasse diagram has edges
$\{(m_{1},e_{1}),(m_{2},e_{1}),(m_{2},e_{2}),(m_{3},e_{2}),(m_{i},e_{3}),(m_{4},e_{3})\};$
see Fig.˜1. Then each inclusion $P_{i}\hookrightarrow Q$ is an initial functor. Note that as posets $P_{1}\cong P_{3}\not\cong P_{2}$.
$m_{1}$$m_{2}$$m_{3}$$m_{4}$$e_{1}$$e_{2}$$e_{3}$
$m_{1}$$m_{2}$$m_{3}$$m_{4}$$e_{1}$$e_{2}$$e_{3}$
Figure 1. The Hasse diagram of the poset $Q$ from Example˜2.13 is shown on the left. The Hasse diagram of each subposet $P_{i}\subseteq Q$ has all of the blue edges and exactly one of the red edges shown on the right.
@@ 2.4. Functors from Posets to Vector Spaces
A generating set of a $Q$-module $G$ is a set $S\subseteq\bigsqcup_{q\in Q}G_{q}$ such that for any $v\in\bigsqcup_{q\in Q}G_{q}$ we have
@@
$\mathbbm{k}^{Q}_{q}$ $=\begin{cases}\mathbbm{k}&\text{if }q\in Q\\ 0&\text{otherwise,}\end{cases}$ $\mathbbm{k}^{Q}_{p,q}=\begin{cases}\mathrm{Id}_{\mathbbm{k}}&{\text{if }p,q\in Q},\\ 0&{\text{otherwise}}.\end{cases}$
@@ 2.4.1. Free Modules
A $Q$-module $G$ is free if there exists a multiset $\mathcal{B}$ of elements in $Q$ such that $G\cong\bigoplus_{q\in\mathcal{B}}\,\mathbbm{k}^{\operatorname{up}[\{q\},Q]}$. The rank of a free $Q$-module $G$ is $|\mathcal{B}|$, i.e., the number of indecomposable summands of $G$. A basis $B$ of a free $Q$-module $G$ is a minimal generating set. We define a function $\beta^{G}\colon Q\to\mathbb{N}$ by $\beta^{G}(q)=|\{s\in B\mid\mathrm{gr}(s)=q\}|$; an elementary linear algebra argument shows that $\beta^{G}$ is independent of the choice of $B$.
Given a morphism $\gamma\colon G\to G^{\prime}$ of finitely generated free $Q$-modules and a choice of ordered bases $B=\{b_{1},\ldots,b_{n}\}$ and $B^{\prime}=\{b^{\prime}_{1},\ldots,b^{\prime}_{m}\}$ for $G$ and $G^{\prime}$, we can represent $G$ as a matrix $\gamma]$, with each row and each column is labeled by an element of $Q$: The matrix is the representation of $\operatorname{colim}\gamma\colon\operatorname{colim}G\to\operatorname{colim}G^{\prime}$ with respect to the ordered bases induced by $B$ and $B^{\prime}$; the $i^{\mathrm{th}}$ column label is $\mathrm{gr}(b_{i})$; and the $i^{\mathrm{th}}$ row label is $\mathrm{gr}(b^{\prime}_{i})$. Regarding $\gamma$ as a functor $Q\times\{0,1\}\to\mathbf{Vec}$, one easily checks that $\gamma$ can be recovered up to isomorphism from $\gamma]$.
@@ 3.1. Structure of Minimal Initial Functors
Our first main result, Theorem˜3.5 below, gives a simple characterization of the minimal initial functors $F\colon P\to Q$ for a large class of posets $Q$. Specifically, the theorem shows that such functors are the same, up to unique isomorphism, as initial scaffolds, which we now define.
@@
For $Q$ a poset with finite downsets, an initial scaffold of $Q$ is a subposet $j\colon P\hookrightarrow Q$ of the following form:
@@
In Example˜2.13, each of the three subposets $P_{1},P_{2},P_{3}\subseteq Q$ is an initial scaffold of $Q$.
@@
It is easily checked that any interval $Q\subseteq\mathbb{N}^{2}$ has a unique initial scaffold $P$, which is the full subposet of $I_{Q}$ in $Q$. Further, $P$ is a zigzag poset, i.e., $P$ is isomorphic to a finite interval of the full subposet of $\mathbb{Z}^{2}$ with the following elements:
$\{(z,z)\mid z\in\mathbb{Z}\}\cup\{(z,z+1)\mid z\in\mathbb{Z}\}.$
@@
    an initial scaffold $j\colon P\hookrightarrow Q$ exists,
@@
    the inclusion $j$ is a minimal initial functor,
@@
    if $Q$ has a finite initial scaffold, then for any minimal initial functor $F\colon C\to Q$, there is an initial scaffold $j\colon P\hookrightarrow Q$ and a unique isomorphism $E\colon C\to P$ such that $F=j\circ E$:
[…]
@@
In October 2025, Brüstle et al. posted a paper giving a variant of Remark˜3.6 under a weaker condition on the poset $Q$ [brustle2025generalized, Theorem C]. Previously, in October 2024, one of us, Lesnick, gave a talk on early versions of our results [Lesnick2024InitialHulls], which centered the result of Remark˜3.6. As one of the authors of [brustle2025generalized] was present for this talk, those authors attribute a version of their Theorem C to us in their paper.
In addition, a variant of our Proposition˜8.2 was obtained independently in [brustle2025generalized]. The paper [brustle2025generalized] also contains other related results which do not overlap with our work. In contrast to our work, [brustle2025generalized] does not deal explicitly with algorithmic questions.
@@ 3.2. Size of Initial Scaffolds
However, in the special case that $Q\subseteq\mathbb{N}^{d}$ is an interval, we show that $|P|$ is controlled by $m$, as follows:
@@
We prove Theorem˜3.10 in Section˜5. The case $d=3$ of Theorem˜3.10 is arguably the most interesting one. As noted in Section˜1, we prove the upper bounds of Theorem˜3.10 by applying an upper bound from [bayer1998monomial] on the size of monomial ideals. The lower bound of Theorem˜3.10 (i) is essentially trivial, while the lower bound of Theorem˜3.10 (ii) is obtained by an explicit construction.
@@
Lemma 3.12 (Dickson’s Lemma [dickson1913finiteness]).
@@ 3.3. Computing Initial Scaffolds
We next state our bounds on the complexity of computing an initial scaffold, deferring the proofs and the algorithms underlying the bounds to Section˜6.
Our first bound is for an arbitrary finite poset:
@@
In the case that $Q\subseteq\mathbb{N}^{d}$ is a finite interval, every vertex is incident to $O(1)$ edges, so the bound of Theorem˜3.13 simplifies to $O(|V|^{2})=O(|\operatorname{Ob}Q|^{2})$. Since $|\operatorname{Ob}Q|$ may be quite large compared to the amount of data required to specify $Q$, one might hope to improve this bound. To this end, for $Q\subseteq\mathbb{N}^{d}$ an interval, we give two specialized algorithms for computing initial scaffolds. Both take as input a different (and usually smaller) representation of $Q$, which we now introduce:
@@
Given an interval $Q\subseteq\mathbb{N}^{d}$, let $Q^{\prime}\coloneqq\operatorname{up}[Q,\mathbb{N}^{d}]\setminus Q$. We call the pair $(M_{Q},M_{Q^{\prime}})$ of sets of minima the upset presentation of $Q$. We let $\|Q\|=|M_{Q}|+|M_{Q^{\prime}}|$.
It is easily checked that an interval $Q\subseteq\mathbb{N}^{d}$ is determined by its upset presentation. Informally, we think of $M_{Q}$ and $M_{Q^{\prime}}$ as the sets of birth and death points of $Q$, respectively. We emphasize that $\|Q\|$ and $|Q|$ are different quantities. Note that by Lemma˜3.12, $\|Q\|$ is always finite.
@@
Remark 3.16.
If $Q$ is a finite interval in $\mathbb{N}^{d}$, then one can instead specify $Q$ by its extrema (i.e., sets of minimal and maximal elements). However, Example˜3.15 (ii) and (iii) illustrate that an infinite interval $Q\subseteq\mathbb{N}^{2}$ is generally not determined by its extrema.
Throughout, we assume that when a finite interval $Q\subset\mathbb{N}^{d}$ is specified by its extrema, the maxima and minima are given separately.
The following result bounds the cost of computing the initial scaffold of an interval in $\mathbb{N}^{d}$.
@@
Given the upset presentation of an interval $Q\subseteq\mathbb{N}^{d}$ with $\|Q\|=s$, we can compute an initial scaffold of $Q$ in time
@@
    $O(s^{4})$ for $d>3$.
In the case that $Q$ is finite, these bounds also hold when $Q$ is instead specified by its extrema, with $s$ the total number of these.
Separate algorithms underlie the bounds (i) and (ii) of Theorem˜3.17. In the case $d=3$, the algorithm for (i) constructs the initial scaffold by iterating through two-dimensional slices of $\mathbb{N}^{3}$. The algorithm for (ii) hinges on the observation that it suffices to restrict attention to the subposet of $Q$ formed by minima and joins of pairs minima.

### v2

@@ 2.2.1. Limits as the Solutions to Linear Systems
For any small category $C$ and functor $G\colon C\to\mathbf{Vec}$, $\lim G$ admits a simple, concrete description as the solution of a system of linear equations: For $v\in\prod_{c\in\operatorname{Ob}C}G_{c}$ and $b\in\operatorname{Ob}C$, let $v_{b}$ be the projection of $v$ onto $G_{b}$. For $\gamma\in\hom C$, write the source and target of $\gamma$ as $s(\gamma)$ and $t(\gamma)$, respectively. A standard result [33, Theorem V.2.2] is that
@@
with each cone map $\lim G\to G_{b}$ the restriction of the projection
$\prod_{c\in\operatorname{Ob}C}G_{c}\twoheadrightarrow G_{b}.$
@@
(2.2) $\lim G=\left\{v\in\!\prod_{m\in M}G_{m}\,:\,G_{lq}(v_{l})=G_{mq}(v_{m})\ \forall\,l,m\in M\textup{ with }l,m<q\right\},$
@@
$\lim G\hookrightarrow\prod_{m\in M}G_{m}\twoheadrightarrow G_{l}\xrightarrow{G_{lq}}G_{q},$
for any choice of $l\in M$ with $l\leq q$.
@@
Sketch of proof of Proposition˜2.10.
@@ 2.3. Initial Functors
  • objects are the pairs $(c,\gamma)$, where $c\in\operatorname{Ob}C$ and $\gamma\in\hom(F_{c},d)$,
  • morphisms $(c,\gamma)\to(c^{\prime},\gamma^{\prime})$ are morphisms $\kappa\colon c\to c^{\prime}$ in $C$ such that $\gamma=\gamma^{\prime}\circ F_{\kappa}$.
@@
The proof of Proposition˜2.14 in [33] gives a concrete construction of the cone $\lim G$ from the cone $\lim(G\circ F)$, where each cone map is taken to be the composition of a cone map of $\lim(G\circ F)$ with a structure map of $G$. This construction makes precise the idea that to compute $\lim G$, it suffices to compute $\lim(G\circ F)$.
We will be particularly interested in functors $F\colon C\to Q$ where $Q$ is a poset. In this case, for $q\in Q$, $(F\downarrow q)$ can be identified with the full subcategory of $C$ with objects $\{c\in\operatorname{Ob}C\mid F_{c}\leq q\}$. In particular, this yields the following observation:
@@
$\{(t,x),(u,x),(u,y),(v,y),(x,z),(y,z),(w,z)\};$
see Fig.˜1(a). For each $i\in\{1,2,3\}$, let $P_{i}\subseteq Q$ be the subposet with $\operatorname{Ob}P_{i}=\operatorname{Ob}Q$, whose Hasse diagram has edges
$\{(t,x),(u,x),(u,y),(v,y),(w,z),e_{i}\},$
where $e_{1}=(t,z)$, $e_{2}=(u,z)$, and $e_{3}=(v,z)$; see Fig.˜1(b). Then each inclusion $P_{i}\hookrightarrow Q$ is an initial functor. Note that as posets $P_{1}\cong P_{3}\not\cong P_{2}$.
ttuuvvwwxxyyzz
(a) The Hasse diagram of the poset $Q$ from Example˜2.17.
ttuuvvwwxxyyzz
(b) The Hasse diagram of each subposet $P_{i}\subseteq Q$ has all of the blue edges and exactly one of the red edges.
Figure 1.
@@ 2.4. Functors from Posets to Vector Spaces
A generating set of a $Q$-module $G$ is a set $S\subseteq\bigsqcup_{q\in Q}G_{q}$ such that for any $v\in\bigsqcup_{q\in Q}G_{q}$, we have
@@
$\mathbbm{k}^{Q}_{q}$ $=\begin{cases}\mathbbm{k}&\text{if }q\in Q,\\ 0&\text{otherwise,}\end{cases}$ $\mathbbm{k}^{Q}_{p,q}=\begin{cases}\mathrm{Id}_{\mathbbm{k}}&{\text{if }p,q\in Q},\\ 0&{\text{otherwise}}.\end{cases}$
@@ 2.4.1. Free Modules
A $Q$-module $G$ is free if there exists a multiset $\mathcal{B}$ of elements in $Q$ such that
$G\cong\bigoplus_{q\in\mathcal{B}}\,\mathbbm{k}^{\operatorname{up}[\{q\},Q]}.$
The rank of a free $Q$-module $G$ is $|\mathcal{B}|$, i.e., the number of indecomposable summands of $G$. A basis $B$ of a free $Q$-module $G$ is a minimal generating set. We define a function $\beta^{G}\colon Q\to\mathbb{N}$ by
$\beta^{G}(q)=|\{s\in B\mid\mathrm{gr}(s)=q\}|,$
where $B$ is any basis of $G$; an elementary linear algebra argument shows that $\beta^{G}$ is independent of the choice of $B$.
Given a morphism $\gamma\colon G\to G^{\prime}$ of finitely generated free $Q$-modules and a choice of ordered bases $B=\{b_{1},\ldots,b_{n}\}$ and $B^{\prime}=\{b^{\prime}_{1},\ldots,b^{\prime}_{m}\}$ for $G$ and $G^{\prime}$, we represent $\gamma$ as a matrix $\gamma]$ with coefficients in $\mathbbm{k}$, with each row and each column labeled by an element of $Q$, as follows:
  • The $j^{\mathrm{th}}$ column of $\gamma]$ is the unique one such that
    $\gamma(b_{j})=\sum_{i:\mathrm{gr}(b_{i}^{\prime})\leq\mathrm{gr}(b_{j})}[\gamma]_{ij}G_{\mathrm{gr}(b_{i}^{\prime})\mathrm{gr}(b_{j})}(b_{i}^{\prime})$
    and $\gamma]_{ij}=0$ if $\mathrm{gr}(b_{i}^{\prime})\not\leq\mathrm{gr}(b_{j})$.
  • The $i^{\mathrm{th}}$ row label is $\mathrm{gr}(b^{\prime}_{i})$.
  • The $j^{\mathrm{th}}$ column label is $\mathrm{gr}(b_{j})$
Regarding $\gamma$ as a functor $Q\times\{0,1\}\to\mathbf{Vec}$, one easily checks that $\gamma$ can be recovered up to natural isomorphism from $\gamma]$.
@@ 3.1. Structure of Minimal Initial Functors
Our first main result, Theorem˜3.5 below, gives a simple characterization of minimal initial functors $F\colon P\to Q$, for a large class of posets $Q$. Specifically, the theorem shows that, up to canonical isomorphism, such functors are the inclusions of certain subposets of $Q$ called initial scaffolds, which we now define:
@@
For $Q$ a poset with finite downsets, an initial scaffold of $Q$ is a subposet $P\subseteq Q$ of the following form:
@@
In Example˜2.17, each of the three subposets $P_{1},P_{2},P_{3}\subseteq Q$ is an initial scaffold of $Q$.
@@
Given an interval $Q\subseteq\mathbb{N}^{2}$, order $M_{Q}$ by $x$-coordinate, writing $M_{Q}=(m_{1},\ldots,m_{k})$. Then
$I_{Q}=M_{Q}\cup\{m_{i}\vee m_{i+1}\mid 1\leq i\leq k-1\},$
where $\vee$ denotes the join operator. The poset $Q$ has a unique initial scaffold $P$, which is the full subposet of $\mathbb{N}^{2}$ with objects $I_{Q}$. Note that $P$ is a zigzag poset; see Definition˜2.5. For instance, if $k=3$, then $P$ is isomorphic to the poset of Example˜2.6.
@@
    an initial scaffold $P\subseteq Q$ exists,
@@
    the inclusion $P\hookrightarrow Q$ is a minimal initial functor,
@@
    if $Q$ has a finite initial scaffold and $F\colon C\to Q$ is a minimal initial functor, then $F$ is an embedding and $\operatorname{im}F$ is an initial scaffold of $Q$.
@@
In October 2025, Brüstle et al. posted a paper giving a variant of Remark˜3.6 under a weaker condition on the poset $Q$ [8, Theorem C]. Previously, in October 2024, one of us, Lesnick, gave a talk on early versions of our results [32], which centered on the result of Remark˜3.6. As one of the authors of [8] was present for this talk, [8] attributes a version of their Theorem C to us.
In addition, a variant of our Proposition˜8.2 was obtained independently in [8]. The paper [8] also contains other related results which do not overlap with our work. In contrast to our work, [8] does not explicitly deal with algorithmic questions.
@@ 3.2. Size of Initial Scaffolds
However, in the special case that $Q\subseteq\mathbb{N}^{d}$ is an interval, we show that $|P|$ is controlled by $n$, as follows:
@@
We prove Theorem˜3.10 in Section˜5. The case $d=3$ of Theorem˜3.10 is arguably the most interesting one. As noted in Section˜1, we prove the upper bounds of Theorem˜3.10 by applying an upper bound from [5] on the size of monomial ideals. The lower bound of Theorem˜3.10 (i) is essentially trivial, while the lower bound of Theorem˜3.10 (ii) is obtained by an explicit construction.
@@
Lemma 3.12 (Dickson’s Lemma [20]).
@@ 3.3. Computing Initial Scaffolds
We next state our bounds on the complexity of computing an initial scaffold, deferring the proofs and the algorithms underlying the bounds to Section˜6. Our first bound is for an arbitrary finite poset:
@@
In the case that $Q\subseteq\mathbb{N}^{d}$ is a finite interval with $d$ constant, every vertex is incident to $O(1)$ edges, so the bound of Theorem˜3.13 simplifies to $O(|V|^{2})=O(|\operatorname{Ob}Q|^{2})$. Since $|\operatorname{Ob}Q|$ may be quite large compared to the amount of data required to specify $Q$, one might hope to improve this bound. To this end, we give two specialized algorithms for computing an initial scaffold of an interval $Q\subseteq\mathbb{N}^{d}$, one for the cases $d=2,3$, and one for arbitrary $d$. Both take as input a different (and usually smaller) representation of $Q$, which we now introduce:
@@
Given an interval $Q\subseteq\mathbb{N}^{d}$, let $Q^{\prime}=\operatorname{up}[Q,\mathbb{N}^{d}]\setminus Q$. We call the pair $(M_{Q},M_{Q^{\prime}})$ of sets of minima the upset presentation of $Q$. We let $\|Q\|=|M_{Q}|+|M_{Q^{\prime}}|$.
It is easily checked that an interval $Q\subseteq\mathbb{N}^{d}$ is determined by its upset presentation. Informally, we think of $M_{Q}$ and $M_{Q^{\prime}}$ as the sets of birth and death points of $Q$, respectively. We emphasize that $\|Q\|$ and $|Q|$ are different quantities, the latter being the number of relations of $Q$, which can be infinite. Note that by Lemma˜3.12, $\|Q\|$ is always finite.
@@
If $Q$ is a finite interval in $\mathbb{N}^{d}$, then one can instead specify $Q$ by its extrema (i.e., sets of minimal and maximal elements). However, Example˜3.15 (ii) and (iii) illustrate that an infinite interval $Q\subseteq\mathbb{N}^{2}$ is generally not determined by its extrema. Throughout, we assume that when a finite interval $Q\subseteq\mathbb{N}^{d}$ is specified by its extrema, the maxima and minima are given separately.
The following result, which we prove in Section˜6, bounds the cost of computing an initial scaffold of an interval in $\mathbb{N}^{d}$.
@@
Given the upset presentation of an interval $Q\subseteq\mathbb{N}^{d}$ with $\|Q\|=n$, we can compute an initial scaffold of $Q$ in time
@@
    $O(n^{4})$ for $d>3$.
Separate algorithms underlie the bounds (i) and (ii) of Theorem˜3.16. In the case $d=3$, the algorithm for (i) constructs the initial scaffold by iterating through two-dimensional slices of $\mathbb{N}^{3}$. The algorithm for (ii) hinges on the observation that it suffices to restrict attention to the subposet of $Q$ formed by minima and joins of pairs of minima.
It is possible that for small enough $d$, say $d=4$, the bound of Theorem˜3.16 (ii) could be improved by extending our algorithm for the case $d=3$ to higher dimensions, but it seems that it would not be trivial to carry out this extension.